OpenStudy (anonymous):

if y= (3)/(4+x^2), then (dy)/(dx)= if y= (3)/(4+x^2), then (dy)/(dx)= @Mathematics

OpenStudy (anonymous):

use the quotient rule

OpenStudy (anonymous):

yeah i know, im confused on the steps of how to do it

OpenStudy (anonymous):

i would not use the quotient rule

OpenStudy (anonymous):

i would use \[(\frac{1}{f})'=-\frac{f'}{f^2}\]

OpenStudy (anonymous):

there is no variable in the numerator

OpenStudy (anonymous):

so answer is \[\frac{-3\times 2x}{(4+x^2)^2}\]

OpenStudy (anonymous):